If it takes 3000 gallons to fill 3/4 of a swimming pool and 2000 gallons to fill a second pool 1/3 full, what percent of the capacity of the first swimming pool is the capacity of the second pool?
Answer by LinnW(1048)

3000/(3/4) = x/1 , 1 represents a full pool
3000*4/3 = x
1000*4 = x
4000 = x
%%%%%%%%%%
2000/(1/3) = y/1
2000*3/1 = y
6000 = y
percent capacity = 6000/4000 = 6/4 = 3/2 = 150%